OpenStudy (anonymous):

Solve the equation for y: 4x - 2y = 6. Answer y = 2x + 3 y = -2x - 3 y = 2x - 3 y = -2x + 3

OpenStudy (anonymous):

Do you know where to start? Which on do you think it is?

OpenStudy (anonymous):

idk how to start

OpenStudy (anonymous):

4x - 2y = 6 -2y = -4x+6 now just divide -2 from -2y = -4x+6 and you will have your answer

OpenStudy (anonymous):

-2x+3?

OpenStudy (anonymous):

2x-3